OpenStudy (anonymous):

RADICAL EXPRESSION 4√125 PLEASE SOLVE NO ITS NOT DIVIDING ITS RADICAL EXPRESSION

OpenStudy (anonymous):

So, you solve the actual answer?

OpenStudy (anonymous):

see 125 =(5*5*5) hence sqrt(125)==sqrt(5*5*5)=5*sqrt(5) thus 4*sqrt(125)=4*5*sqrt(5) (usesqrt(5)=2.236 (aapx) =20*2.236 ==44.721

OpenStudy (anonymous):

no decimals

OpenStudy (anonymous):

because i cant submit my problem like that

OpenStudy (anonymous):

then what u want ot oevaluate (u mau leave it ilike 20sqrt(5)

OpenStudy (anonymous):

well yes i have that answer thank u